OpenStudy (anonymous):

derivative of sqrt ((x^2 - 9)/(x+3))

OpenStudy (anonymous):

\[\sqrt ((x^2 - 9)/(x+3))\]

OpenStudy (anonymous):

with steps please

OpenStudy (anonymous):

radical over numerator or whole thing?

OpenStudy (anonymous):

whole thing

OpenStudy (anonymous):

then make your life easier and write \[\sqrt{\frac{x^2-9}{x+3}}=\sqrt{\frac{(x+3)(x-3)}{x+3}}=\sqrt{x-3}\]

OpenStudy (anonymous):

OMG didn't even think of that lol

OpenStudy (anonymous):

could you help with \[f(z)= 1/(e^z +1)^2\]